OpenStudy (anonymous):

find f(3) if f(x)=3^3+2x^2-x-1. If possible please show me how to solve this.

OpenStudy (texaschic101):

basically, you are just going to sub in 3 everywhere there is an x. f(3) = 3^3 + 2(3^2) - 3 - 1 f(3) = 27 + 2(9) - 4 f(3) = 27 + 18 - 4 f(3) = 45 - 4 f(3) = 41

OpenStudy (anonymous):

okay, thanks. Math is not my best subject.